Construct an example to show that the cross product of two vectors in 3 d does not transform like a vector under a reflection and therefore is a pseudo vector

1
Expert's answer
2021-12-06T16:42:51-0500

A pseudovector is a quantity that is similar to a vector but undergoes an additional inversion under a coordinate reflection.

For example, "x\times y=z"

Under a reflection in the xy plane, this equation becomes "x\times y=-z"


example:

"x=(1,2,3),y=(0,1,2)"


"x\times y=\begin{vmatrix}
 i & j&k \\
 1 & 2&3\\
0&1&2
\end{vmatrix}=i-2j+k"


"y\times x=\begin{vmatrix}
 i & j&k \\
 0 & 1&2\\
1&2&3
\end{vmatrix}=-i+2j-k=-(x\times y)"